# Frustration-free free fermions

### v1

@@ Abstract
We develop a general theory of frustration-free free-fermion systems and derive their necessary and sufficient conditions. Assuming locality and translation invariance, we find that the possible band touching between the valence bands and the conduction bands is always quadratic or softer, which rules out the possibility of describing Dirac and Weyl semimetals using frustration-free local Hamiltonians. We further construct a frustration-free free-fermion model on the honeycomb lattice and show that its density fluctuations acquire an anomalous gap originating from the diverging quantum metric associated with the quadratic band-touching points. Nevertheless, an $O(1/L^{2})$ finite-size scaling of the charge-neutral excitation gap can be verified even in the presence of interactions, consistent with the more general results we derive in an accompany work [arXiv:2503.12879].
@@
More technically, we say a Hamiltonian is frustration-free if all the local terms are simultaneously minimized by the ground states [1]. Frustration-freeness serves as a powerful starting point for deriving general results about both the ground state and the excitation spectrum of the system—a rare possibility for quantum many-body problems. Restrictive as it may seem, a surprisingly diverse array of quantum phases could be realized using frustration-free Hamiltonains. Many important spin models, such as the Affleck–Kennedy–Lieb–Tasaki model [2, 1], the Majumdar–Ghosh (MG) model [3, 4], and more generally the parent Hamiltonians of matrix product states [5, 6] are all frustration-free.
Paradoxical as it may seem, frustration-free models also serve as important theoretical tools for investigating exotic quantum phases associated with geometric frustration. Examples include the Rokhsar-Kivelson model [7] for quantum critical points and the Kitaev toric code models [8] for quantum spin liquids, as well as the free-fermion nearest-neighbor hopping models on the kagome, pychlore, and other line-graph lattices [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. The frustration-freeness also allows one to derive rigorous results about the nature of the ground state(s) when electron-electron interaction is incorporated [9, 10, 12, 14, 11, 13, 21, 22, 23], as exemplified by the saturated ferromagnetism in partially filled flat bands [1]. These results are particularly relevant in recent years due to the emergence of tunable flat-band materials in both natural [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36] and artificial [37, 38, 39, 40, 41, 42] platforms. Existing general results on frustration-free Hamiltonians typically concern bosonic systems like quantum magnets [43, 44, 45, 46, 47, 48, 49, 50]. Motivated by the recently highlighted connections between frustration-freeness, flat bands, and correlated topological materials [1, 51, 52], we develop a general theory for frustration-free free fermions (F4) in this work.
Frustration-free conditions—We begin by deriving the F4 conditions for particle-number conserving Hamiltonians. Consider a free-fermion Hamiltonian $\hat{H}$ defined on a lattice
@@
Here, $\hat{H}_{\bm{R}}$ is a local term defined around the point $\bm{R}$, $\hat{\bm{c}}_{\bm{R}}$ is an $N_{r}$-dimensional column vector with components being the $\hat{c}_{\bm{r}\sigma}$ within a range $|\bm{r}-\bm{R}|<r$, and $\sigma$ labels the fermion modes (say sublattices, orbitals, and spin) assigned to $\bm{r}$. We assume $\hat{H}$ is local in the sense that $r$ is finite and independent of the system size. The fermion hopping amplitudes are encoded in the $N_{r}$-dimensional Hermitian matrices $h_{\bm{R}}$. The constant $C_{\bm{R}}$ will be fixed later.
Let $\mu_{\bm{R}\alpha}>0$ ($\alpha=1,2,\cdots,A_{\bm{R}}$) be positive eigenvalues of $h_{\bm{R}}$ and $\bm{\psi}_{\bm{R}\alpha}$ be the corresponding normalized eigenvectors. Likewise, let $-\nu_{\bm{R}\beta}<0$ ($\beta=1,2,\cdots,B_{\bm{R}}$) and $\bm{\phi}_{\bm{R}\beta}$ be the negative eigenvalues and associated eigenvectors. $h_{\bm{R}}$ might admit null vectors but they will not be used. We define corresponding annihilation operators by
@@
Fermions with the same label $\bm{R}$ satisfy canonical anticommutation relations due to the orthonormality of the eigenvectors. Also, as both $\hat{\psi}$ and $\hat{\phi}$ are sums of fermion annihilation operators, for any $\bm{R},\bm{R}^{\prime}$ we have
@@
For $\bm{R}\neq\bm{R}^{\prime}$, however, $\{\hat{\psi}_{\bm{R}\alpha},\hat{\psi}_{\bm{R}^{\prime}\alpha^{\prime}}^{\dagger}\}$, $\{\hat{\phi}_{\bm{R}\beta},\hat{\phi}_{\bm{R}^{\prime}\beta^{\prime}}^{\dagger}\}$, and $\{\hat{\psi}_{\bm{R}\alpha},\hat{\phi}_{\bm{R}^{\prime}\beta}^{\dagger}\}$ for $\bm{R}\neq\bm{R}^{\prime}$ do not vanish generally.
@@
Note that both $\hat{\psi}_{\bm{R}\alpha}^{\dagger}\hat{\psi}_{\bm{R}\alpha}$ and $\hat{\phi}_{\bm{R}\beta}\hat{\phi}_{\bm{R}\beta}^{\dagger}=1-\hat{\phi}_{\bm{R}\beta}^{\dagger}\hat{\phi}_{\bm{R}\beta}$ are projectors with eigenvalues 00 and $1$. Therefore, $\hat{H}_{\bm{R}}$ is positive semi-definite. As the choice of the constants $C_{\bm{R}}$ is immaterial, our discussion up to this point applies generally to any number-conserving free-fermion problems.
@@
We now introduce our first F4 result. Recall a local Hamiltonian $\hat{H}$ is frustration-free if there exists a split $\hat{H}=\sum_{\bm{R}}\hat{H}_{\bm{R}}$ with positive-semidefinite local terms $\hat{H}_{\bm{R}}$ and a state $|\Phi\rangle$ such that $\hat{H}_{\bm{R}}|\Phi\rangle=0$ for all $\bm{R}$ simultaneously [1]. We claim a free-fermion $\hat{H}$ expressed as in Eq. (4) is frustration-free if and only if
@@
$\{\hat{\psi}_{\bm{R}\alpha},\hat{\phi}_{\bm{R}^{\prime}\beta}^{\dagger}\}=0\quad\text{for all $\bm{R}$, $\bm{R}^{\prime}$, $\alpha$, $\beta$.}$ (5)
@@
To appreciate the validity of the condition, we note that if $\hat{H}$ is F4 then its ground state $|\Phi\rangle$ satisfies $\hat{\psi}_{\bm{R}\alpha}|\Phi\rangle=\hat{\phi}_{\bm{R}^{\prime}\beta}^{\dagger}|\Phi\rangle=0$ for all $\bm{R},\alpha,\beta$. As $\{\hat{\psi}_{\bm{R}\alpha},\hat{\phi}_{\bm{R}^{\prime}\beta}^{\dagger}\}$ is a $\mathbb{C}$-number, the existence of $|\Phi\rangle$ implies the anti-commutator vanishes. In the other direction, one can first define a state $|\Phi\rangle$ by applying all of the linearly independent fermion creation operators among the set $\{\hat{\phi}_{\bm{R},\beta}^{\dagger}\}_{\bm{R},\beta}$ to the vacuum [53]. The vanishing of the anti-commutator in Eq. (5) implies $|\Phi\rangle$ is annihilated by all $\hat{H}_{\rm R}$, proving $\hat{H}$ is F4.
@@
We note that the condition Eq. (5) also extends naturally to fermionic bilinear Hamiltonians without particle U(1) symmetry, as we discuss in Ref. [53].
Dirac trick—It is generally known that Dirac-like Hamiltonians on a bipartite lattice are related to frustration-free models defined on its sublattices. For instance, the F4 model on the checkerboard lattice can be obtained by squaring the single-particle Hamiltonian associated with the Lieb lattice [54, 9, 10]. We now show that similar relations can be obtained for any F4 models.
From the split in Eq. (4), we can write the full Hamiltonian as $\hat{H}=\hat{H}^{(+)}+\hat{H}^{(-)}$, where $\hat{H}^{(\pm)}=\sum_{\bm{R}}\hat{H}^{(\pm)}_{\bm{R}}$. By construction, the non-zero single-particle energies $\omega^{(+)}_{\gamma}$ associated with $\hat{H}^{(+)}$ (resp. $\omega^{(-)}_{\gamma}$ for $\hat{H}^{(-)}$) are all positive (negative). Furthermore, if $\hat{H}$ is F4, Eq. (5) implies $[\hat{H}_{\bm{R}}^{(+)},\hat{H}_{\bm{R}^{\prime}}^{(-)}]=0$ and therefore the non-zero single-particle energies of $\hat{H}$ can be obtained by combining those from $\hat{H}^{(\pm)}$.
We now show that the non-zero single-particle energies can always be written as $\omega^{(\pm)}_{\gamma}=\pm\left(\nu^{(\pm)}_{\gamma}\right)^{2}$, where $\nu^{(\pm)}$ are obtained from the associated Dirac-like Hamiltonian. For concreteness, we focus on $\hat{H}^{(+)}$ below and the construction for $\hat{H}^{(-)}$ is similar. From the form in Eq. (4), we can construct an auxiliary Hamiltonian $\hat{\tilde{H}}^{(+)}=\sum_{\bm{R}}\hat{\tilde{H}}_{\bm{R}}^{(+)}$, where
$\begin{split}\hat{\tilde{H}}_{\bm{R}}^{(+)}=\left(\begin{array}[]{cc}\hat{\bm{c}}_{\bm{R}}^{\dagger}&\hat{\bm{d}}_{\bm{R}}^{\dagger}\end{array}\right)\left(\begin{array}[]{cc}0&\bm{\psi}_{\bm{R}}\sqrt{\mu_{\bm{R}}}\\ \sqrt{\mu_{\bm{R}}}\bm{\psi}_{\bm{R}}^{\dagger}&0\end{array}\right)\left(\begin{array}[]{c}\hat{\bm{c}}_{\bm{R}}\\ \hat{\bm{d}}_{\bm{R}}\end{array}\right).\end{split}$ (6)
Here, $\hat{\bm{d}}_{\bm{R}}=(\hat{d}_{\bm{R}1},\hat{d}_{\bm{R}2},\dots,\hat{d}_{\bm{R}A_{\bm{R}}})^{T}$ is a vector of auxiliary fermion annihilation operators introduced at $\bm{R}$, $\mu_{\bm{R}}$ is a diagonal matrix consisting of the original local eigenvalues $\mu_{\bm{R}\alpha}>0$ for $\alpha=1,\dots,A_{\bm{R}}$, and $\psi_{\bm{R}}$ is an $N_{r}\times A_{\bm{R}}$ dimensional matrix formed by the eigenvectors $\bm{\psi}_{\bm{R}\alpha}$ of $h_{\bm{R}}$. As defined, $\hat{\tilde{H}}^{(+)}$ is a local Hamiltonian with chiral symmetry, and so its non-zero single-particle energies come in pairs of the form ±$\pm\nu^{(+)}_{\gamma}$. Although the “square root” is taken for each local term $\hat{H}_{\bm{R}}^{(+)}$, one can verify that the decoupled block in $\left(\tilde{h}^{(+)}\right)^{2}$ acting on the $\hat{c}$-fermions is identical to the single-particle Hamiltonian $h^{(+)}$ of the original problem (Appendix A). This establishes $\omega^{(+)}=\left(\nu^{(+)}_{\gamma}\right)^{2}$.
@@
Excitation gap—Our next result concerns the finite-size scaling of the excitation gap of F4 Hamiltonians. Consider a family of F4 Hamiltonians $\{\hat{H}(L)\}$ labeled by the finite linear size $L$. Let $\epsilon(L)>0$ be the first non-zero eigenvalue of $\hat{H}(L)$. We say the system is gapless if $\epsilon(L)$ approaches 00, the ground-state energy, in the thermodynamic limit. In the following, we use the Dirac trick to prove that $\epsilon(L)\sim O(1/L^{2})$ for gapless F4 Hamiltonians with particle U(1) and lattice translation symmetries.
From the F4 condition, $[\hat{H}^{(+)},\hat{H}^{(-)}]=0$ and so the Bloch Hamiltonians associated with $\hat{H}^{(+)}$ and $\hat{H}^{(-)}$ commute at every momentum $\bm{k}$ in the Brillouin zone (BZ). The mentioned non-zero single-particle energies of $\hat{H}$ can then be labeled by the momentum $\bm{k}$ as $\omega^{(\pm)}_{\bm{k},n}$, where $n=1,2,\dots$ is counted from zero energy. Furthermore, the auxiliary Dirac-like Hamiltonians are also translation-invariant, and we have the momentum-resolved relation $\omega^{(\pm)}_{\bm{k},n}=\pm\left(\nu_{\bm{k},n}^{(\pm)}\right)^{2}$. This reveals the non-zero bands of any F4 model can always be understood as the square of some auxiliary bands.
If $\hat{H}$ is gapless, at least one of $\nu_{\bm{k},1}^{(+)}$ and $\nu_{\bm{k},1}^{(-)}$ must vanish at some momentum $\bm{k}_{*}$. Consider a small line segment $\bm{k}=\bm{k}_{*}+\delta k\hat{\bm{\kappa}}$ passing through $\bm{k}_{*}$ for any direction $\hat{\bm{\kappa}}$. As the auxiliary Hamiltonians only contain finite-range hopping, with suitable labeling the dispersion $\nu^{(\pm)}_{\bm{k},n}$ can also be expanded analytically in $\delta k$ [55]. Now, if $\nu_{\bm{k}_{*},1}^{(+)}=0$, we have $\nu^{(+)}_{\bm{k},1}=v_{\hat{\bm{\kappa}}}\delta k+O(\delta k^{2})$ for some velocity component $v_{\hat{\bm{\kappa}}}$. We can then conclude a quadratic dispersion $\omega_{\bm{k}1}^{(+)}=v_{\hat{\bm{\kappa}}}^{2}\delta k^{2}+\mathcal{O}(\delta k^{3})$ around $\bm{k}=\bm{k}_{0}+\delta k\hat{\bm{\kappa}}$. The $O(1/L^{2})$ finite-size scaling of excitation energy then follows from the discrete sampling of the BZ.
We also comment briefly on the case without translation symmetry. A main difficulty in this case is the lack of a simple way to define a sequence of Hamiltonians $\{\hat{H}(L)\}$ approaching the thermodynamic limit. Yet, if we assume the weaker condition $\epsilon(L)\sim O(1/L)$ for local free-fermion Hamiltonians, then the stronger conclusion $\epsilon(L)\sim O(1/L^{2})$ for F4 Hamiltonians follows from the Dirac trick.
Honeycomb model—We now illustrate our results through an example. Consider spinless fermions hopping on the two-dimensional honeycomb lattice with one $s$ orbital per site (Fig. 1). As is realized in graphene, there are symmetry-protected band touchings at the K and K’ points. Although linear dispersion is generally expected, our results imply any F4 Hamiltonians on the honeycomb must have the Dirac fermions softened into band touching points that are quadratic or softer. In other words, the F4 requirement automatically tunes the Dirac velocity to zero, as we show below in a concrete model.
Let $\hat{c}^{\dagger}_{{\bm{R}},\sigma}$ for $\sigma=A,B$ be spinless fermions defined on the $\sigma$ site of unit cell ${\bm{R}}$. We consider a finite $L$-by-$L$ lattice along the two lattice vectors. Based on the F4 condition Eq. (5), we define compactly supported fermions $\hat{\psi}_{\bm{R}}$ and $\hat{\phi}_{\bm{R}}$ by taking superposition of the six orbitals around a hexagon, as shown in Fig. 1 (a,b). One can verify $\{\hat{\psi}_{\bm{R}},\hat{\phi}_{\bm{R}^{\prime}}^{\dagger}\}=0$ and therefore the Hamiltonain $\hat{H}=t\sum_{\bm{R}}\left(\hat{\psi}_{\bm{R}}^{\dagger}\hat{\psi}_{\bm{R}}+\hat{\phi}_{\bm{R}}\hat{\phi}_{\bm{R}}^{\dagger}\right)$ is F4.
Upon Fourier transform with $\hat{c}_{{\bm{k}},\sigma}^{\dagger}=\frac{1}{L}\sum_{{\bm{R}}}\hat{c}_{{\bm{R}},\sigma}^{\dagger}e^{i{\bm{k}}\cdot{\bm{R}}}$, one finds
@@
$\begin{split}\hat{\psi}_{{\bm{k}}}^{\dagger}=\hat{\bm{c}}^{\dagger}_{{\bm{k}}}{\bm{\psi}}_{{\bm{k}}};\quad{\bm{\psi}}_{{\bm{k}}}=\frac{1}{\sqrt{6}}\left(\begin{array}[]{c}1+e^{i2\pi k_{2}}+e^{i2\pi(k_{1}+k_{2})}\\ 1+e^{i2\pi k_{1}}+e^{i2\pi(k_{1}+k_{2})}\end{array}\right),\end{split}$ (7)
@@
where we write ${\bm{k}}=k_{1}{\bm{b}}_{1}+k_{2}{\bm{b}}_{2}$ with ${\bm{b}}_{1,2}$ denoting the reciprocal lattice vectors and $\hat{\bm{c}}^{\dagger}_{{\bm{k}}}=(\hat{c}_{{\bm{k}},A}^{\dagger},\hat{c}_{{\bm{k}},b}^{\dagger})$. $\hat{\phi}_{{\bm{k}}}^{\dagger}$ is similarly defined but with ${\bm{\phi}}_{{\bm{k}}}=\sigma_{3}{\bm{\psi}}_{{\bm{k}}}$. Here, $\sigma_{3}$ is the Pauli matrix in the sublattice space. Since the fermions defined for different unit cells do not generally anti-commute with each other, ${\bm{\psi}}_{{\bm{k}}}$ and ${\bm{\phi}}_{{\bm{k}}}$ as defined are not normalized. Instead, one has $|{\bm{\psi}}_{{\bm{k}}}|^{2}=|{\bm{\phi}}_{{\bm{k}}}|^{2}=e_{\bm{k}}$, where

### v2

@@ Abstract
We develop a general theory of frustration-free free-fermion systems and derive the necessary and sufficient conditions for such Hamiltonians. Assuming locality and translation invariance, we find that any band touching between the valence and the conduction bands is always quadratic or softer, which rules out the possibility of describing Dirac and Weyl semimetals using frustration-free local Hamiltonians. We further construct a frustration-free free-fermion model on the honeycomb lattice and show that its density fluctuations acquire an anomalous gap originating from the diverging quantum metric associated with the quadratic band-touching points. Nevertheless, an $O(1/L^{2})$ finite-size scaling of the charge-neutral excitation gap can be verified even in the presence of interactions, consistent with the more general results we derive in an accompanying work [arXiv:2503.12879].
@@
More technically, we say a Hamiltonian is frustration-free if all the local terms are simultaneously minimized by the ground states [1]. Frustration-freeness serves as a powerful starting point for deriving general results about both the ground state and the excitation spectrum of the system—a rare possibility for quantum many-body problems. Restrictive as it may seem, a surprisingly diverse array of quantum phases can be realized using frustration-free Hamiltonians. Many important spin models are frustration-free, such as the Affleck–Kennedy–Lieb–Tasaki model [2, 1], the Majumdar–Ghosh (MG) model [3, 4], and more generally the parent Hamiltonians of matrix product states [5, 6].
Paradoxical as it may seem, frustration-free models also serve as important theoretical tools for investigating exotic quantum phases associated with geometric frustration. Examples include the Rokhsar-Kivelson model [7] for quantum critical points and the Kitaev toric code models [8] for quantum spin liquids, as well as the free-fermion nearest-neighbor hopping models on the kagome, pyrochlore, and other line-graph lattices [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Frustration-freeness also allows one to derive rigorous results about the nature of the ground state(s) when electron-electron interactions are incorporated [9, 10, 12, 14, 11, 13, 21, 22, 23], as exemplified by the saturated ferromagnetism in partially filled flat bands [1]. These results are particularly relevant in recent years due to the emergence of tunable flat-band materials in both natural [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36] and artificial [37, 38, 39, 40, 41, 42] platforms. Existing general results on frustration-free Hamiltonians typically concern bosonic systems like quantum magnets [43, 44, 45, 46, 47, 48, 49, 50]. Motivated by the recently highlighted connections between frustration-freeness, flat bands, and correlated topological materials [1, 51, 52], we develop a general theory for frustration-free free fermions (F4) in this work.
F4 conditions—We begin by deriving the F4 conditions for particle-number conserving Hamiltonians. Consider a free-fermion Hamiltonian $\hat{H}$ defined on a lattice
@@
Here, $\hat{H}_{\bm{R}}$ is a local term defined around the point $\bm{R}$, $\hat{\bm{c}}_{\bm{R}}$ is an $N_{r}$-dimensional column vector with components given by $\hat{c}_{\bm{r}\sigma}$ within a range $|\bm{r}-\bm{R}|<r$, and $\sigma$ labels the fermion modes (e.g., sublattices, orbitals, and spin) assigned to $\bm{r}$. We assume that $\hat{H}$ is local in the sense that $r$ is finite and independent of the system size. The fermion hopping amplitudes are encoded in the $N_{r}$-dimensional Hermitian matrices $h_{\bm{R}}$. The constant $C_{\bm{R}}$ will be fixed later.
Let $\mu_{\bm{R}\alpha}>0$ ($\alpha=1,2,\cdots,A_{\bm{R}}$) be the positive eigenvalues of $h_{\bm{R}}$ and $\bm{\psi}_{\bm{R}\alpha}$ be the corresponding normalized eigenvectors. Likewise, let $\nu_{\bm{R}\beta}<0$ ($\beta=1,2,\cdots,B_{\bm{R}}$) and $\bm{\phi}_{\bm{R}\beta}$ be the negative eigenvalues and associated eigenvectors. We do not use the null vectors in our construction. Define
@@
Operators with the same label $\bm{R}$ automatically satisfy canonical anticommutation relations. Also, for any $\bm{R},\bm{R}^{\prime}$
@@
For $\bm{R}\neq\bm{R}^{\prime}$, however, $\{\hat{\psi}_{\bm{R}\alpha},\hat{\psi}_{\bm{R}^{\prime}\alpha^{\prime}}^{\dagger}\}$, $\{\hat{\phi}_{\bm{R}\beta},\hat{\phi}_{\bm{R}^{\prime}\beta^{\prime}}^{\dagger}\}$, and $\{\hat{\psi}_{\bm{R}\alpha},\hat{\phi}_{\bm{R}^{\prime}\beta}^{\dagger}\}$ do not vanish generally.
@@
Note that both $\hat{\psi}_{\bm{R}\alpha}^{\dagger}\hat{\psi}_{\bm{R}\alpha}$ and $\hat{\phi}_{\bm{R}\beta}\hat{\phi}_{\bm{R}\beta}^{\dagger}=1-\hat{\phi}_{\bm{R}\beta}^{\dagger}\hat{\phi}_{\bm{R}\beta}$ are projectors with eigenvalues 0 and $1$. Therefore, $\hat{H}_{\bm{R}}$ is positive semi-definite.
@@
Our discussion up to this point applies generally to any number-conserving free-fermion problems. We now introduce our first result—the necessary and sufficient condition for F4 models. Recall that a local Hamiltonian $\hat{H}$ is frustration-free if it admits a decomposition $\hat{H}=\sum_{\bm{R}}\hat{H}_{\bm{R}}$ with positive-semidefinite local terms $\hat{H}_{\bm{R}}$ and a state $|\Phi\rangle$ such that $\hat{H}_{\bm{R}}|\Phi\rangle=0$ for all $\bm{R}$ simultaneously [1]. Given the form in Eq. (4), we claim
@@
$\hat{H}~\text{is F${}^{4}$}\quad\Longleftrightarrow\quad\{\hat{\psi}_{\bm{R}\alpha},\hat{\phi}_{\bm{R}^{\prime}\beta}^{\dagger}\}=0\quad\forall\bm{R},\bm{R}^{\prime},\alpha,\beta.$ (5)
@@
A sketch of the proof is provided below; a more detailed discussion can be found in Ref. [53]. If $\hat{H}$ is F4, then its ground state $|\Phi\rangle$ satisfies $\hat{\psi}_{\bm{R}\alpha}|\Phi\rangle=\hat{\phi}_{\bm{R}^{\prime}\beta}^{\dagger}|\Phi\rangle=0$ for all $\bm{R},\alpha,\beta$. As $\{\hat{\psi}_{\bm{R}\alpha},\hat{\phi}_{\bm{R}^{\prime}\beta}^{\dagger}\}$ is a $\mathbb{C}$-number, the existence of $|\Phi\rangle$ implies the anti-commutator vanishes. In the other direction, one can define a state $|\Phi\rangle$ by applying all of the linearly independent fermion creation operators among the set $\{\hat{\phi}_{\bm{R},\beta}^{\dagger}\}_{\bm{R},\beta}$ to the vacuum. The vanishing of the anti-commutator in Eq. (5) then implies $\hat{H}_{\bm{R}}|\Phi\rangle=0$ for all $\bm{R}$.
@@
We note that the condition Eq. (5) also generalizes naturally to fermionic bilinear Hamiltonians without particle U(1) symmetry, which, as we discuss in Ref. [53], goes beyond some existing results [54, 55].
@@
Excitation gap of gapless models—As a first application of the F4 condition, we prove that the excitation gap $\epsilon(L)$ scales as $O(1/L^{2})$ for gapless F4 Hamiltonians with particle U(1) and lattice translation symmetries, where $L$ denotes the finite linear size of the system.
More concretely, consider a F4 Hamiltonian $\hat{H}(L)$ with a variable finite linear size $L$. Let $\epsilon(L)>0$ be the first non-zero eigenvalue of $\hat{H}(L)$ (possibly in a different fermion number sector). We say the system is gapless if $\epsilon(L)$ approaches 0, the ground-state energy, in the thermodynamic limit. From the decomposition in Eq. (4), we can write the full Hamiltonian as $\hat{H}=\hat{H}^{(+)}+\hat{H}^{(-)}$, where $\hat{H}^{(\pm)}=\sum_{\bm{R}}\hat{H}^{(\pm)}_{\bm{R}}$. By construction, the non-zero single-particle energies $\omega^{(+)}_{\gamma}$ associated with $\hat{H}^{(+)}$ (resp. $\omega^{(-)}_{\gamma}$ for $\hat{H}^{(-)}$) are all positive (negative). Furthermore, the F4 condition Eq. (5) implies $\hat{H}_{\bm{R}}^{(+)},\hat{H}_{\bm{R}^{\prime}}^{(-)}]=0$ and therefore the non-zero single-particle energies of $\hat{H}$ can be obtained by combining those from $\hat{H}^{(\pm)}$. Upon Fourier transform, the Bloch Hamiltonians associated with $\hat{H}^{(+)}$ and $\hat{H}^{(-)}$ commute at every momentum $\bm{k}$ in the Brillouin zone (BZ). The mentioned non-zero single-particle energies of $\hat{H}$ can then be labeled by the momentum $\bm{k}$ as $\omega^{(\pm)}_{\bm{k},n}$, where $n=1,2,\dots$ is counted from zero energy.
As $\omega^{(\pm)}_{\bm{k},n}$ are independently defined from $\hat{H}^{(\pm)}$, one heuristically expects any band touching between the two sets, if any, must be quadratic or softer. The finite-size sampling on the dispersion curve then leads to the asserted finite-size gap scaling. This heuristic can be justified more carefully through the construction of an auxiliary Dirac Hamiltonian. For concreteness, we focus on $\hat{H}^{(+)}$ below; the construction for $\hat{H}^{(-)}$ is similar. Starting from Eq. (4), we can construct an auxiliary Hamiltonian $\hat{\tilde{H}}^{(+)}=\sum_{\bm{R}}\hat{\tilde{H}}_{\bm{R}}^{(+)}$, where
@@
$\begin{split}\hat{\tilde{H}}_{\bm{R}}^{(+)}=\left(\begin{array}[]{cc}\hat{\bm{c}}_{\bm{R}}^{\dagger}&\hat{\bm{d}}_{\bm{R}}^{\dagger}\end{array}\right)\left(\begin{array}[]{cc}0&{\psi}_{\bm{R}}\sqrt{\mu_{\bm{R}}}\\ \sqrt{\mu_{\bm{R}}}{\psi}_{\bm{R}}^{\dagger}&0\end{array}\right)\left(\begin{array}[]{c}\hat{\bm{c}}_{\bm{R}}\\ \hat{\bm{d}}_{\bm{R}}\end{array}\right).\end{split}$ (6)
Here, $\hat{\bm{d}}_{\bm{R}}=(\hat{d}_{\bm{R}1},\hat{d}_{\bm{R}2},\dots,\hat{d}_{\bm{R}A_{\bm{R}}})^{T}$ is a vector of auxiliary fermion annihilation operators introduced at $\bm{R}$, $\mu_{\bm{R}}$ is a diagonal matrix consisting of the original local eigenvalues $\mu_{\bm{R}\alpha}>0$ for $\alpha=1,\dots,A_{\bm{R}}$, and $\psi_{\bm{R}}$ is an $N_{r}\times A_{\bm{R}}$ dimensional matrix formed by the eigenvectors $\bm{\psi}_{\bm{R}\alpha}$ of $h_{\bm{R}}$.
As defined, $\hat{\tilde{H}}^{(+)}$ is a local Hamiltonian with chiral symmetry, and so its non-zero single-particle energies come in pairs of the form $\pm\nu^{(+)}_{\gamma}$. Although the “square root” is taken for each local term $\hat{H}_{\bm{R}}^{(+)}$, one can verify that the decoupled block in $\left(\tilde{h}^{(+)}\right)^{2}$ acting on the $\hat{c}-$fermions is identical to the single-particle Hamiltonian $h^{(+)}$ of the original problem (Appendix A). This establishes $\omega^{(+)}=\left(\nu^{(+)}_{\gamma}\right)^{2}$. Furthermore, the auxiliary Dirac-like Hamiltonians are also translation-invariant, and we have the momentum-resolved relation $\omega^{(\pm)}_{\bm{k},n}=\pm\left(\nu_{\bm{k},n}^{(\pm)}\right)^{2}$. This reveals that the non-zero bands of any F4 model can always be understood as the square of some auxiliary bands.
If $\hat{H}$ is gapless, at least one of $\nu_{\bm{k},1}^{(+)}$ and $\nu_{\bm{k},1}^{(-)}$ must vanish at some momentum $\bm{k}_{*}$. Consider a small line segment $\bm{k}=\bm{k}_{*}+\delta k\hat{\bm{\kappa}}$ passing through $\bm{k}_{*}$ for any direction $\hat{\bm{\kappa}}$. As the auxiliary Hamiltonians only contain finite-range hopping, with suitable labeling the dispersion $\nu^{(\pm)}_{\bm{k},n}$ can also be expanded analytically in $\delta k$ [56]. Now, if $\nu_{\bm{k}_{*},1}^{(+)}=0$, we have $\nu^{(+)}_{\bm{k},1}=v_{\hat{\bm{\kappa}}}\delta k+O(\delta k^{2})$ for some velocity component $v_{\hat{\bm{\kappa}}}$. We can then conclude a quadratic dispersion $\omega_{\bm{k}1}^{(+)}=v_{\hat{\bm{\kappa}}}^{2}\delta k^{2}+\mathcal{O}(\delta k^{3})$ around $\bm{k}=\bm{k}_{*}+\delta k\hat{\bm{\kappa}}$. The $O(1/L^{2})$ finite-size scaling of excitation energy follows from the discrete sampling of the BZ.
Honeycomb model—We now illustrate our results through an example. Consider spinless fermions hopping on the two-dimensional honeycomb lattice with one $s$ orbital per site (Fig. 1). There are symmetry-protected band touchings at the K and K’ points. Although linear dispersion, like that is graphene, is generally expected, our results imply that any F4 Hamiltonians must have the Dirac fermions softened into band touching points that are quadratic or softer, i.e., the F4 requirement automatically tunes the Dirac velocity to zero.
Let $\hat{c}^{\dagger}_{{\bm{R}},\sigma}$ for $\sigma=A,B$ be spinless fermions defined on the $\sigma$ site of unit cell $\bm{R}$}. We consider a finite $L$-by-$L$ lattice along the two lattice vectors. Based on the F4 condition Eq. (5), we define compactly supported fermions $\hat{\psi}_{\bm{R}}$ and $\hat{\phi}_{\bm{R}}$ by taking a superposition of the six orbitals around a hexagon, as shown in Fig. 1 (a,b). One can verify $\{\hat{\psi}_{\bm{R}},\hat{\phi}_{\bm{R}^{\prime}}^{\dagger}\}=0$ and therefore $\hat{H}=t\sum_{\bm{R}}\left(\hat{\psi}_{\bm{R}}^{\dagger}\hat{\psi}_{\bm{R}}+\hat{\phi}_{\bm{R}}\hat{\phi}_{\bm{R}}^{\dagger}\right)$ is a F4 Hamiltonian.
Introduce Fourier transform with $\hat{c}_{{\bm{k}},\sigma}^{\dagger}=\frac{1}{L}\sum_{{\bm{R}}}\hat{c}_{{\bm{R}},\sigma}^{\dagger}e^{i{\bm{k}}\cdot{\bm{R}}}$:
$\begin{split}\hat{\psi}_{{\bm{k}}}^{\dagger}=\hat{\bm{c}}^{\dagger}_{{\bm{k}}}{\bm{\psi}}_{{\bm{k}}};\quad{\bm{\psi}}_{{\bm{k}}}=\frac{1}{\sqrt{6}}\left(\begin{array}[]{c}1+e^{i2\pi k_{2}}+e^{i2\pi(k_{1}+k_{2})}\\ 1+e^{i2\pi k_{1}}+e^{i2\pi(k_{1}+k_{2})}\end{array}\right)\end{split}$ (7)
@@
where we write ${\bm{k}}=k_{1}{\bm{b}}_{1}+k_{2}{\bm{b}}_{2}$ with ${\bm{b}}_{1,2}$ denoting the reciprocal lattice vectors and $\hat{\bm{c}}^{\dagger}_{{\bm{k}}}=(\hat{c}_{{\bm{k}},A}^{\dagger},\hat{c}_{{\bm{k}},b}^{\dagger})$. $\hat{\phi}_{{\bm{k}}}^{\dagger}$ is similarly defined but with ${\bm{\phi}}_{{\bm{k}}}=\sigma_{3}{\bm{\psi}}_{{\bm{k}}}$. Here, $\sigma_{3}$ is the Pauli matrix in the sublattice space. Since the fermions defined on neighboring unit cells do not satisfy canonical anti-commutation relations, ${\bm{\psi}}_{{\bm{k}}}$ and ${\bm{\phi}}_{{\bm{k}}}$ as defined are not normalized. Instead, one has $|{\bm{\psi}}_{{\bm{k}}}|^{2}=|{\bm{\phi}}_{{\bm{k}}}|^{2}=e_{\bm{k}}$, where